# The SHMS 11 GeV/c Spectrometer in Hall C at Jefferson Lab

### v1

@@ 1.1 Jefferson Lab Overview
The Continuous Electron Beam Accelerator Facility at Thomas Jefferson National Accelerator Facility (Jefferson Lab) provides high energy electron beams for fundamental nuclear physics experiments. Originally planned for maximum electron beam energies of 4 GeV, the accelerator operated at energies of up to 6 GeV starting in 2000. An upgrade of the facility was completed in 2017, enabling beam delivery at a maximum energy of 12 GeV to the new experimental Hall D, and 11 GeV to the existing Halls, A, B, and C [1]. (Fig. 1.)
@@ 1.2 Hall C Experimental Program at 6 GeV
In addition, the relatively small minimum angle of 10.5 degrees accessible with the HMS allowed the execution of pion electroproduction experiments where, in many cases, the pion is emitted in the forward direction. This allowed the successful execution of a program of measurements of the pion form factor [3, 4], which also incorporated precise L-T separations, as well measurements of charged pion production in Semi-inclusive Deep Inelastic Scattering (SIDIS) [5] (see Figs. 4 and 5).
The momentum reach of the HMS, up to the maximum beam energy of 6 GeV, enabled measurements of the process to large [6, 7]. This allowed a search for signs of color transparency (Fig. 6) as well as measurements of inclusive electron scattering at to access contributions of “superfast” quarks to inelastic structure functions [8] and measure the relative contributions of Short Range Correlations (SRCs) in the nuclear wave function [9] (Fig. 7).
@@ 1.3 Hall C 12 GeV Program
The first commissioning experiment was a measurement of inclusive electron scattering cross sections from hydrogen and deuterium [11] (see Fig. 8). Such a cross section experiment is an excellent testing ground for understanding of the spectrometer acceptance, while not pushing the SHMS performance in other areas. Some settings for this experiment were chosen to allow simultaneous measurement with the well-understood HMS to provide a cross check. In addition, some time was devoted to the measurement of inclusive cross section ratios for nuclear targets relative to deuterium [12]. These ratios are well-measured for certain nuclei and serve as another straightforward verification of the spectrometer acceptance due to the need to compare yields from extended (10 cm long) targets to shorter, solid targets (mm scale). These measurements resulted in the first extraction of the EMC Effect in 10B and 11B [13].
@@
An extension of the 6 GeV color transparency experiments to larger [14] served as an excellent first experiment with which to exercise the SHMS in coincidence mode. In this experiment, there are few random coincidences, so isolating the coincidence reaction is straightforward. This experiment, as well as a measurement of deuteron electro-disintegration [15], also tested the high momentum capabilities of the SHMS, exceeding GeV/ in these experiments. Although the maximum central momentum of the SHMS is almost 11 GeV, the momentum of GeV/ was already sufficient to learn about the performance of the superconducting magnets and spectrometer optics when pushed to a significant fraction of the spectrometer’s ultimate capabilities. In addition, the body of elastic data acquired from both these initial coincidence experiments provided constraints on the experiment kinematics, testing the possible variation of, e.g. the spectrometer pointing or central momentum for various settings. Results from the color transparency and deuteron electro-disintegration experiments are shown in Figs. 9 and 10.
@@
A set of meson electroproduction experiments followed the initial commissioning experiments and further exercised the SHMS capabilities. Two of the experiments measured charged pion electroproduction in semi-inclusive deep inelastic scattering, SIDIS [18, 19]. The SHMS was used at central angles smaller than 7∘ for the SIDIS running, a fact which contributed to relatively high singles rates in the SHMS. Both experiments aimed to make precise measurements of ratios, so control of rate dependent systematic effects was a key challenge. The third experiment [20] measured exclusive cross sections for production above the resonance region, extracting the longitudinal and transverse cross sections via a Rosenbluth separation. In this case, the experimental uncertainties were expected to be dominated by statistics rather than systematics, so this served as an excellent candidate for a first L-T separation. In common with the charged pion SIDIS experiments, the kaon experiment required use of the SHMS at small angles and so faced high singles rates.
@@
The “commissioning” and “year-1” experiments described above give a sense of the SHMS capabilities important for the overall physics program. Since then, a variety of experiments have been completed in Hall C. These include measurements of photoproduction [21], virtual Compton scattering [22], exclusive charged pion electroproduction to extract the pion form factor and for cross section scaling tests [23], inclusive electron scattering from polarized 3He to extract and [24, 25], and exclusive and inclusive scattering from nuclei to make measurements of short range correlations and the EMC Effect [26, 27, 12]. In the future, additional L-T separations in inclusive scattering (to measure from hydrogen, deuterium, and several nuclei) and semi-inclusive reactions (to make the first precise measurement of for the SIDIS reaction) are also planned. While not all future Hall C experiments will make use of the SHMS, it is a key component of its 12 GeV experimental program.
@@ 2 Specifications for the upgraded Hall-C Spectrometer complex
| Target Length Accepted at 90∘(HMS)/45∘ (SHMS) | 10 cm | 25 cm |
@@
| e/h Discrimination | 1000:1 at | |
| 98% efficiency | 1000:1 at | |
| 98% efficiency | ||
| /K Discrimination | 100:1 at | |
| 95% efficiency | 100:1 at | |
| 95% efficiency |
@@ 3.1.1 The Magnets and Vacuum Channel
As shown in Fig. 11, in order to fit within the space available in Hall C, the SHMS must be even shorter than its lower-momentum partner, the HMS. All of the SHMS magnets are superconducting so that they can provide the necessary large bending and focusing effects in the relatively short available distance. Given the requirement to access small scattering angles, the HB and the first two quadrupoles (Q1 and Q2) have special provisions to provide clearance for the electron beam and its vacuum pipe. HB is a “C”-magnet so that all of the flux-return iron is on the side away from the beamline. (See Fig. 12.) As initially constructed, the HB leaked significant field into the beamline such that the beam would have been deflected outside of the beam dump. Simulations were done to determine the optimal shielding design to reduce the field in the beamline region for all combinations of SHMS angle and momentum and these mitigations were implemented [28]. The front of the HB cryostat, between the beamline and the magnet bore, is made very narrow. Both Q1 and Q2 have notches in their cryostats and iron yokes so that they, too, can clear the beamline when the spectrometer is configured at small scattering angles. Yoke steel for Q1 is inside the cryostat. The final quadrupole (Q3) and the dipole (DSHMS) have external warm yokes. Parameters of the SHMS magnets are provided in Table 2.
To minimize multiple scattering as particles pass through the SHMS, the bores of all of the magnets are evacuated. The vacuum space begins at a window on the front of HB. The entrance window into the HB is approximately 15 cm square and is made of 0.01” thick aluminum. A vacuum connection is made between the exit of HB and Q1 entrance which is followed by the 40 cm diameter vacuum bore in Q1. The exit of Q1 is connected to the entrance of Q2 by a vacuum pipe. The vacuum vessel bore through Q2, Q3, and DSHMS is 60 cm in diameter. The location of the end of the vacuum after the exit of DSHMS depends on the needs of the experiment. If the experiment needs the Noble Gas Cherenkov (NGC) detector (described in Sec. 4.5), then a window is placed at the exit of DSHMS with the NGC detector placed between the exit window and the drift chambers. Otherwise, a Vacuum Extension Tank (VET) is attached to the exit of the DSHMS that puts the exit window at 30 cm from the first drift chamber in the detector stack. In both cases, the dipole exit window is made of 0.020” thick aluminum.
@@ 3.1.2 Charged Particle Transport Models
A magnetic transport code, SNAKE [29], was used to model the transport of charged particles in the SHMS. The SNAKE model of the SHMS incorporated the mechanical sizes of the magnets, while the magnetic fields were generated by the static field analysis code TOSCA [30] and validated with field measurements. The relative strengths of the field integrals of the magnets were selected to maximize the acceptance while simultaneously providing the desired momentum and scattering angle resolutions. For a charged particle with relative momentum, , where is the momentum of the particle and is the central momentum of the spectrometer, the transport from the target to the focal plane located midway between the two drift chambers can be expressed in terms of a matrix representation of the solutions of the equation of motion of charged particles in magnetic fields [31]. The first-order transport matrix for the SHMS is given by:
@@
The acceptance of the spectrometer is mainly determined by the collimator placed within a remotely operated collimator box that is that installed between the HB magnet and the first quadrupole magnet. The collimator ladder assembly within this box may be positioned at one of three settings. The top position (accessed when the assembly is at its lowest position) is a stretched octagon with an opening height 9.843 in and width 6.693 in on the upstream side. It is 2.5 in thick. The lower two positions contain sieve slits with holes in a rectangular pattern separated by 0.6457 in horizontally and 0.9843 in vertically. The sieve pattern at the middle ladder position has 11 columns of holes with the sixth column centered horizontally. The holes on the bottom sieve are in ten columns and are offset by one half column gap from those in the middle sieve. The sieve collimators are 1.25 in thick. The geometry is illustrated in Fig. 13. Both sieves and the octagonal collimator are made of Mi-TechTM Tungsten HD-17 (Density 17 g/cm3, and composition 90% W, 6% Ni, 4% Cu) [32].
@@
The SNAKE-based transport model of the SHMS was also used to study the spectrometer acceptance. The acceptance of the SHMS as a function of , as determined by the SNAKE based model is plotted in Fig. 15. Alternatively, the spectrometer acceptance is also modeled using the Hall C Monte Carlo (SIMC) with the magnetic transport matrix obtained from the COSY INFINITY [33] program. The acceptance of the SHMS as a function of as determined by SIMC is also plotted in Fig. 15. The two acceptance models agree within statistics.
@@ 3.2 Shield House Layout, Shielding Design
The radiation environment is an important consideration for the design of the SHMS shield house, in particular the effect of radiation-induced effects on the performance and reliability of detectors and electronics. It has been shown that many new commercial off the shelf components are sensitive to radiation damage and single event upsets, requiring a careful evaluation of the impact of the radiation-induced effects on their performance and reliability [34, 35]. A specialized SHMS shield house design was thus developed. Shielding thicknesses were optimized using a Monte Carlo simulation and benchmarked against the HMS shielding house, which has been proven to provide the necessary detector shielding over more than a decade of experiments at the 6 GeV JLab. A full description of the shielding optimization can be found in Ref. [36].
@@
Primary and scattered electrons lose a significant amount of energy as they traverse a material by producing a large number of lower energy photons through bremsstrahlung [37]. It is thus important to consider shielding materials that efficiently stop the latter as well.
@@
Typical beam-target geometries were simulated using Monte Carlo techniques. Simulations were performed using the GEANT MCWORKS distribution, which includes detailed physical and geometric descriptions of the experimental hall and simulates the physics processes using standard GEANT3 together with the DINREG nuclear fragmentation package. Hadronic interactions are treated using the DINREG package, which calculates the probability of such interactions using a database of photonuclear cross sections. For electron-nucleus interactions, an “equivalent photon” representation of the electron (or positron) is used.
@@
A limitation of the radiation studies is the lack of cross section data for low-energy neutrons. The accuracy of the GEANT simulations was tested by benchmark calculations using the MCNP code [38] with an isotropic neutron point source of 1 MeV located 1 m from the shield wall. The MCNP calculations suggest that 50 cm of concrete thermalizes most of the fast neutrons, and after 1 m practically no epithermal neutrons remain. The thermalized neutrons can be captured by a 1 cm Boron layer. In reality, however, the neutron spectrum also includes higher energy neutrons, for instance produced by electrons interacting in the concrete, and thus the actual amount of material for the walls exposed to the primary sources of radiation has to be thicker. A simple transmission calculation using GEANT4 for incident neutron beams of energies between 1 and 10 MeV suggests that a thickness 150 cm of concrete is sufficient to stop the majority of low-energy neutrons.
The SHMS shielding model is composed of standard concrete ( = 2.4 g/cm-1). The thickness of the wall in front of the detector and electronics rooms is 200 cm, to shield from the primary radiation source around the target. Figure 19 shows the surviving background flux for varying front wall concrete thicknesses. The results are normalized to the background flux in the HMS at 20∘. This angle was chosen as experiments in Hall C have shown that electronics problems are more frequent at lower angles. The simulation results suggest that 200 cm of concrete reduces the total flux to half that of the HMS at 20∘.
@@
The majority of charged particles are stopped by the outer walls of the spectrometer shield house. An additional source of radiation may be created from particles entering the enclosure through the magnets. To protect the electronics further, an intermediate wall was installed between the detector and electronics rooms. Figure 21 shows the normalized rate as the thickness of this intermediate wall is varied. This suggests that the optimal configuration is provided by a concrete thickness of 80–100 cm111Note that a minimum wall thickness of 50 cm is needed to provide support for the roof of the shield house. Further details on shielding configurations investigated and their optimization can be found in Ref. [36].
@@
The incoming photon flux has two components: externally produced photons and bremsstrahlung photons produced by electrons in the twenty radiation lengths of concrete. The simulations have shown that the outgoing photon spectrum is soft (10 MeV). Placing a lead layer after the concrete is essential to suppress this low energy photon flux. The () reaction in lead is not a problem. The threshold for the reaction is given by the neutron binding energy ( 8 MeV). At higher energies, the cross sections are in the mbarn range [39]. Even disregarding the low cross section, however, it is not clear that this reaction contributes to the irradiation of the electronics, because a high energy photon is replaced by a low energy (but not thermal) neutron.
@@
Fig. 22 is a photograph showing the resulting multi-layered shielding in one of the SHMS shield house walls. The ceiling, floor, and other walls have similar compositions but varying dimensions presented earlier in Fig. 18. Details about the development of custom concrete material containing boron can be found in Ref. [40].
In summary, the SHMS shielding consists of concrete walls to moderate and attenuate particles. Low energy (thermal) neutrons are absorbed in a boron layer inside the concrete. Low energy and 0.5 MeV capture photons are absorbed in lead. With this design, the rates at forward angles of 5.5∘ are estimated to be less than 70% of the design goal (HMS at 20∘) in the detector room and below 50% in the electronics room.
4 Design, Development and Calibration of the SHMS Detectors
Basic trigger information comes from four planes of scintillator or fused silica hodoscopes. Tracking is provided by twelve planes of conventional drift chambers. Particle identification uses gas and aerogel Cherenkov counters, a preshower counter, and a total-absorption shower counter.
4.1 Scintillator Trigger Hodoscopes
The SHMS hodoscope system provides a clean trigger and trigger time information as well as the definition of the detector package fiducial area, required for physics cross section measurements. The system is composed of four separate planes of detector paddles: S1X and S1Y located immediately after the second drift chamber, and S2X and S2Y approximately 2.6 m away along the z direction. The S1X, S1Y, and S2X planes were built using thin plastic scintillator paddles while S2Y uses fused silica bars.

### v2

@@ 1.1 Jefferson Lab Overview
The Continuous Electron Beam Accelerator Facility at Thomas Jefferson National Accelerator Facility (Jefferson Lab) provides high energy electron beams for fundamental nuclear physics experiments. Originally planned for maximum electron beam energies of 4 GeV, the accelerator operated at energies of up to 6 GeV starting in 2000. An upgrade of the facility was completed in 2017, enabling beam delivery at a maximum energy of 12 GeV to the new experimental Hall D, and 11 GeV to the existing Halls, A, B, and C [PhysRevAccelBeams.27.084802]. (Fig. 1.)
@@ 1.2 Hall C Experimental Program at 6 GeV
In addition, the relatively small minimum angle of 10.5 degrees accessible with the HMS allowed the execution of pion electroproduction experiments where, in many cases, the pion is emitted in the forward direction. This allowed the successful execution of a program of measurements of the pion form factor [Volmer:2000ek, Horn:2006tm], which also incorporated precise L-T separations, as well measurements of charged pion production in Semi-inclusive Deep Inelastic Scattering (SIDIS) [Navasardyan:2006gv] (see Figs. 4 and 5).
The momentum reach of the HMS, up to the maximum beam energy of 6 GeV, enabled measurements of the process to large [Abbott:1997bc, Garrow:2001di]. This allowed a search for signs of color transparency (Fig. 6) as well as measurements of inclusive electron scattering at to access contributions of “superfast” quarks to inelastic structure functions [Fomin:2010ei] and measure the relative contributions of Short Range Correlations (SRCs) in the nuclear wave function [Fomin:2011ng] (Fig. 7).
@@ 1.3 Hall C 12 GeV Program
The first commissioning experiment was a measurement of inclusive electron scattering cross sections from hydrogen and deuterium [E12-10-002] (see Fig. 8). Such a cross section experiment is an excellent testing ground for understanding of the spectrometer acceptance, while not pushing the SHMS performance in other areas. Some settings for this experiment were chosen to allow simultaneous measurement with the well-understood HMS to provide a cross check. In addition, some time was devoted to the measurement of inclusive cross section ratios for nuclear targets relative to deuterium [E12-10-008]. These ratios are well-measured for certain nuclei and serve as another straightforward verification of the spectrometer acceptance due to the need to compare yields from extended (10 cm long) targets to shorter, solid targets (mm scale). These measurements resulted in the first extraction of the EMC Effect in 10B and 11B [HallC:2022utd].
@@
An extension of the 6 GeV color transparency experiments to larger [E12-06-107] served as an excellent first experiment with which to exercise the SHMS in coincidence mode. In this experiment, there are few random coincidences, so isolating the coincidence reaction is straightforward. This experiment, as well as a measurement of deuteron electro-disintegration [E12-10-003], also tested the high momentum capabilities of the SHMS, exceeding GeV/ in these experiments. Although the maximum central momentum of the SHMS is almost 11 GeV, the momentum of GeV/ was already sufficient to learn about the performance of the superconducting magnets and spectrometer optics when pushed to a significant fraction of the spectrometer’s ultimate capabilities. In addition, the body of elastic data acquired from both these initial coincidence experiments provided constraints on the experiment kinematics, testing the possible variation of, e.g. the spectrometer pointing or central momentum for various settings. Results from the color transparency and deuteron electro-disintegration experiments are shown in Figs. 9 and 10.
@@
A set of meson electroproduction experiments followed the initial commissioning experiments and further exercised the SHMS capabilities. Two of the experiments measured charged pion electroproduction in semi-inclusive deep inelastic scattering, SIDIS [E12-09-017, E12-09-002]. The SHMS was used at central angles smaller than 7∘ for the SIDIS running, a fact which contributed to relatively high singles rates in the SHMS. Both experiments aimed to make precise measurements of ratios, so control of rate dependent systematic effects was a key challenge. The third experiment [E12-09-011] measured exclusive cross sections for production above the resonance region, extracting the longitudinal and transverse cross sections via a Rosenbluth separation. In this case, the experimental uncertainties were expected to be dominated by statistics rather than systematics, so this served as an excellent candidate for a first L-T separation. In common with the charged pion SIDIS experiments, the kaon experiment required use of the SHMS at small angles and so faced high singles rates.
@@
The “commissioning” and “year-1” experiments described above give a sense of the SHMS capabilities important for the overall physics program. Since then, a variety of experiments have been completed in Hall C. These include measurements of photoproduction [Duran:2022xag], virtual Compton scattering [Li:2022sqg], exclusive charged pion electroproduction to extract the pion form factor and for cross section scaling tests [E12-19-006], inclusive electron scattering from polarized 3He to extract and [E12-06-110, E12-06-121], and exclusive and inclusive scattering from nuclei to make measurements of short range correlations and the EMC Effect [E12-17-005, E12-06-105, E12-10-008]. In the future, additional L-T separations in inclusive scattering (to measure from hydrogen, deuterium, and several nuclei) and semi-inclusive reactions (to make the first precise measurement of for the SIDIS reaction) are also planned. While not all future Hall C experiments will make use of the SHMS, it is a key component of its 12 GeV experimental program.
@@ 2 Specifications for the upgraded Hall-C Spectrometer complex
| Target Length Accepted at 90∘(HMS)/40∘ (SHMS) | 10 cm | 25 cm |
@@
| e/h Discrimination | 1000:1 at 98% efficiency | 1000:1 at 98% efficiency |
| /K Discrimination | 100:1 at 95% efficiency | 100:1 at 95% efficiency |
@@ 3.1.1 The Magnets and Vacuum Channel
As shown in Fig. 11, in order to fit within the space available in Hall C, the SHMS must be even shorter than its lower-momentum partner, the HMS. A schematic of the SHMS carriage with the five magnets is shown in Fig. 12. All of the SHMS magnets are superconducting so that they can provide the necessary large bending and focusing effects in the relatively short available distance. Given the requirement to access small scattering angles, the HB and the first two quadrupoles (Q1 and Q2) have special provisions to provide clearance for the electron beam and its vacuum pipe. HB is a “C”-magnet so that all of the flux-return iron is on the side away from the beamline. As initially constructed, the HB leaked significant field into the beamline such that the beam would have been deflected outside of the beam dump. Simulations were done to determine the optimal shielding design to reduce the field in the beamline region for all combinations of SHMS angle and momentum and these mitigations were implemented [Moore_2014]. The front of the HB cryostat, between the beamline and the magnet bore, is made very narrow. Both Q1 and Q2 have notches in their cryostats and iron yokes so that they, too, can clear the beamline when the spectrometer is configured at small scattering angles. Yoke steel for Q1 is inside the cryostat. The final quadrupole (Q3) and the dipole (DSHMS) have external warm yokes. Parameters of the SHMS magnets are provided in Table 2.
To minimize multiple scattering as particles pass through the SHMS, the bores of all of the magnets are evacuated. The vacuum space begins at a window on the front of HB. The entrance window into the HB is approximately 15 cm square and is made of 0.01” thick aluminum. A vacuum connection is made between the exit of HB and Q1 entrance which is followed by the 40 cm diameter vacuum bore in Q1. The exit of Q1 is connected to the entrance of Q2 by a vacuum pipe. The vacuum vessel bore through Q2, Q3, and DSHMS is 60 cm in diameter. The location of the end of the vacuum after the exit of DSHMS depends on the needs of the experiment. If the experiment needs the Noble Gas Cherenkov (NGC) detector (described in Sec. 4.5), then a window is placed at the exit of DSHMS with the NGC detector placed between the exit window and the drift chambers. Otherwise, a Vacuum Extension Tank (VET) is attached to the exit of the DSHMS that puts the exit window at 30 cm from the first drift chamber in the detector stack. In both cases, the dipole exit window is made of 0.020” thick aluminum. The arrangement of the detectors and the distances between the detectors does not change when the NGC is replaced by the VET. All detectors have a fixed location in the SHMS hut on a carriage. Only the aerogel detector (see Fig. 38) is on rails so that it can be easily removed without changing the location of the other detectors.
@@ 3.1.2 Charged Particle Transport Models
A magnetic transport code, SNAKE [snake_code], was used to model the transport of charged particles in the SHMS. The SNAKE model of the SHMS incorporated the mechanical sizes of the magnets, while the magnetic fields were generated by the static field analysis code TOSCA [TOSCA] and validated with field measurements. The relative strengths of the field integrals of the magnets were selected to maximize the acceptance while simultaneously providing the desired momentum and scattering angle resolutions. For a charged particle with relative momentum, , where is the momentum of the particle and is the central momentum of the spectrometer, the transport from the target to the focal plane located midway between the two drift chambers can be expressed in terms of a matrix representation of the solutions of the equation of motion of charged particles in magnetic fields [Penner61]. The first-order transport matrix for the SHMS is given by:
@@
The acceptance of the spectrometer is mainly determined by the collimator placed within a remotely operated collimator box that is installed between the HB magnet and the first quadrupole magnet. The collimator ladder assembly within this box may be positioned at one of three settings. The top position (accessed when the assembly is at its lowest position) is a stretched octagon with an opening height 9.843 in and width 6.693 in on the upstream side. It is 2.5 in thick. The lower two positions contain sieve slits with holes in a rectangular pattern separated by 0.6457 in horizontally and 0.9843 in vertically. The sieve pattern at the middle ladder position has 11 columns of holes with the sixth column centered horizontally. The holes on the bottom sieve are in ten columns and are offset by one half column gap from those in the middle sieve. The sieve collimators are 1.25 in thick. The geometry is illustrated in Fig. 13. Both sieves and the octagonal collimator are made of Mi-TechTM Tungsten HD-17 (Density 17 g/cm3, and composition 90% W, 6% Ni, 4% Cu) [Mitech].
@@
The SNAKE-based transport model of the SHMS was also used to study the spectrometer acceptance. The acceptance of the SHMS as a function of , as determined by the SNAKE based model is plotted in Fig. 15. Alternatively, the spectrometer acceptance is also modeled using the Hall C Monte Carlo (SIMC) with the magnetic transport matrix obtained from the COSY INFINITY [COSY] program. The acceptance of the SHMS as a function of as determined by SIMC is also plotted in Fig. 15. The two acceptance models agree within statistics.
@@ 3.2 Shield House Layout, Shielding Design
The radiation environment is an important consideration for the design of the SHMS shield house, in particular the effect of radiation-induced effects on the performance and reliability of detectors and electronics. It has been shown that many new commercial off the shelf components are sensitive to radiation damage and single event upsets, requiring a careful evaluation of the impact of the radiation-induced effects on their performance and reliability [griff97, under98]. A specialized SHMS shield house design was thus developed. Shielding thicknesses were optimized using a Monte Carlo simulation and benchmarked against the HMS shielding house, which has been proven to provide the necessary detector shielding over more than a decade of experiments at the 6 GeV JLab. A full description of the shielding optimization can be found in Ref. [horn08].
@@
Primary and scattered electrons lose a significant amount of energy as they traverse a material by producing a large number of lower energy photons through bremsstrahlung [leo87]. It is thus important to consider shielding materials that efficiently stop the latter as well.
@@
Typical beam-target geometries were simulated using Monte Carlo techniques. Simulations were performed using the GEANT MCWORKS distribution, which includes detailed physical and geometric descriptions of the experimental hall and simulates the physics processes using standard GEANT3 [GEANT3] together with the DINREG nuclear fragmentation package. Hadronic interactions are treated using the DINREG package, which calculates the probability of such interactions using a database of photonuclear cross sections. For electron-nucleus interactions, an “equivalent photon” representation of the electron (or positron) is used.
@@
A limitation of the radiation studies is the lack of cross section data for low-energy neutrons. The accuracy of the GEANT simulations was tested by benchmark calculations using the MCNP code [bries93] with an isotropic neutron point source of 1 MeV located 1 m from the shield wall. The MCNP calculations suggest that 50 cm of concrete thermalizes most of the fast neutrons, and after 1 m practically no epithermal neutrons remain. The thermalized neutrons can be captured by a 1 cm Boron layer. In reality, however, the neutron spectrum also includes higher energy neutrons, for instance produced by electrons interacting in the concrete, and thus the actual amount of material for the walls exposed to the primary sources of radiation has to be thicker. A simple transmission calculation using GEANT4 [geant4] for incident neutron beams of energies between 1 and 10 MeV suggests that a thickness 150 cm of concrete is sufficient to stop the majority of low-energy neutrons.
The SHMS shielding model is composed of standard concrete ( = 2.4 g/cm-1). The thickness of the wall in front of the detector and electronics rooms is 100 cm, to shield from the primary radiation source around the target. Figure 19 shows the surviving background flux for varying front wall concrete thicknesses. The results are normalized to the background flux in the HMS at 20∘. This angle was chosen as experiments in Hall C have shown that electronics problems are more frequent at lower angles. The simulation results show that 100 cm of concrete reduces the total flux to same as the flux in HMS at 20∘.
@@
The majority of charged particles are stopped by the outer walls of the spectrometer shield house. An additional source of radiation may be created from particles entering the enclosure through the magnets. To protect the electronics further, an intermediate wall was installed between the detector and electronics rooms. Figure 21 shows the normalized rate as the thickness of this intermediate wall is varied. This suggests that the optimal configuration is provided by a concrete thickness of 80–100 cm . Due to space constraints the minimum wall thickness of 50 cm needed to provide support for the roof of the shield house was used. Further details on shielding configurations investigated and their optimization can be found in Ref. [horn08].
@@
The incoming photon flux has two components: externally produced photons and bremsstrahlung photons produced by electrons in the twenty radiation lengths of concrete. The simulations have shown that the outgoing photon spectrum is soft (10 MeV). Placing a lead layer after the concrete is essential to suppress this low energy photon flux. The () reaction in lead is not a problem. The threshold for the reaction is given by the neutron binding energy ( 8 MeV). At higher energies, the cross sections are in the mbarn range [biren95]. Even disregarding the low cross section, however, it is not clear that this reaction contributes to the irradiation of the electronics, because a high energy photon is replaced by a low energy (but not thermal) neutron.
@@
Fig. 22 is a photograph showing the resulting multi-layered shielding in one of the SHMS shield house walls. The ceiling, floor, and other walls have similar compositions but varying dimensions presented earlier in Fig. 18. Details about the development of custom concrete material containing boron can be found in Ref. [concrete].
In summary, the SHMS shielding consists of concrete walls to moderate and attenuate particles. Low energy (thermal) neutrons are absorbed in a boron layer inside the concrete. Low energy and 0.5 MeV capture photons are absorbed in lead. With this design, the rates at forward angles of 5.5∘ are estimated to be less than 70% of the design goal (HMS at 20∘) in the detector room and below 70% in the electronics room.
4 Design, Construction and Calibration of the SHMS Detectors
The layout of the SHMS detectors in the SHMS detector hut is shown in Fig. 23 and each detector is described in detail in the following subsections. A CAD drawing of the SHMS detector stack is shown in Fig. 49. After traveling from the target through the SHMS
magnets, the scattered particles exit the SHMS dipole vacuum pipe. The first detector that is encountered is the Noble Gas Cherenkov detector (NGC) which is described in Sec. 4.5. The NGC can be replaced by a vacuum tank which is designed to occupy the same space as the NGC. The tracking of the particles is done by a pair of drift chambers (DC1 and DC2) and each of drift chambers consists of six wire planes. Details on DC1 and DC2 are described in Sec. 4.3. Basic trigger information comes from four planes of scintillator or fused silica hodoscopes. This first pair of hodoscope planes (S1X and S1Y) are located directly after drift chambers. The Heavy Gas Cherenkov (HGC) detector is placed after S1Y and the HGC is described in Sec. 4.4. The aerogel cherenkov (AER) detector is located after the HGC and is described in Sec. 4.6. The AER can be removed from the detector stack if it is desired without changing the positions of the other detectors. The second pair of hodoscope planes (S2X and S2Y) are located after the aerogel cherenkov detector. The S1X, S1Y and S2X hodoscope planes are standard scintillator paddles and are described in Sec. 4.1. The S2Y hodoscope plane is made of fused silica (”quartz”) and is described in Sec. 4.2. The last detectors in the stack are the preshower and shower detectors which are described in Sec. 4.7.
4.1 Scintillator Trigger Hodoscope
The SHMS hodoscope system provides a clean trigger and trigger time information as well as the definition of the detector package fiducial area, required for physics cross section measurements. The system is composed of four separate planes of detector paddles: S1X and S1Y located immediately after the second drift chamber, and S2X and S2Y approximately 2.6 m away along the z direction. Teh S1X plane is closest to the drift chamber and the S2Y plane is 9.6 cm away from the S1X plane along the z-direction. The S1X, S1Y, and S2X planes were built using thin plastic scintillator paddles while S2Y uses fused silica bars.